# On the Effectiveness of ASR Representations in Real-world Noisy Speech Emotion Recognition

### v3

@@ 2.2 IEMOCAP-Human
To replicate human speech noise in real-world environments, we merged the IEMOCAP data with the Office category from the DEMAND corpus. We specifically used the OMEETING segment of the Office data, recorded in a busy office cafeteria, to augment the IEMOCAP dataset.
@@ IV-B Experimental Procedure
In this study, we conducted six experiments to investigate various aspects of ASR representations in NSER.
In Experiment 1, we examined the impact of ASR representations on NSER across the MELD, IEMOCAP, IEMOCAP-Environment, and IEMOCAP-Human datasets at an SNR of 0 dB. We explored two main strategies for NSER: noise reduction methods and fine-tuning of SSL models. For noise reduction, we employed the speech separation method (Conv-TasNet) and speech enhancement method (DCCRN). In the SSL fine-tuning category, we fine-tuned models such as Wav2vec 2.0111https://huggingface.co/facebook/wav2vec2-base, HuBERT222https://huggingface.co/facebook/hubert-base-ls960, and WavLM333https://huggingface.co/microsoft/wavlm-base on noisy speech data, which had seen widespread use in tasks involving both SER and NSER [56, 57]. Furthermore, we compared the performance of conventional ASR models with large-scale ASR models in NSER tasks. In the conventional ASR, we evaluated a Conformer-based ASR model, which had been extensively utilized in tasks combining emotion recognition and ASR [58, 59, 60]. For large-scale ASR, we selected Whisper owing to its robust multi-language support, multi-task processing capabilities, strong generalization from large-scale pre-training, and robustness in complex acoustic environments.
In Experiment 2, we examined the effectiveness of different ASR representations on NSER across the MELD, IEMOCAP, IEMOCAP-Environment, and IEMOCAP-Human datasets at an SNR of 0 dB. We performed a comparative analysis of the ASR encoder and ASR decoder, evaluating the performance of models using the last layer and the mean of all layer representations. Additionally, we proposed the layer adapter method to fuse representations from different levels, obtaining a more comprehensive ASR representation for NSER.
@@
In Experiment 3, we examined the robustness of ASR representations in NSER under various noise intensities. A comparative analysis was carried out using the IEMOCAP-Environment and IEMOCAP-Human datasets, systematically controlling noise at SNR levels of -5, 0, and 5 dB. We compared different ASR representations, including the encoder with the adapter, the decoder with the adapter, and a combination of the encoder and decoder with the adapter, alongside representations from SSL models.
@@
In Experiment 4, we examined the correlation between ASR performance and NSER outcomes under various noise intensities. A comparative analysis was conducted using the MELD, IEMOCAP, IEMOCAP-Environment, and IEMOCAP-Human datasets. We evaluated ASR representations from two models: a Conformer-based ASR model, representing the conventional ASR approach, and Whisper, representing the large-scale ASR approach.
@@
In Experiment 5, we examined the effectiveness of ASR representations across different modalities in NSER at an SNR of 0 dB. A comparative analysis was carried out between speech- and text-based modalities, leveraging the SSL method on both ASR transcriptions and ground-truth transcriptions. For the text modality, we employed a BERT model, following prior studies on emotion recognition [61, 62]. For the speech modality, we compared different ASR representations from Whisper, including the encoder with the adapter, the decoder with the adapter, and a combination of the encoder and decoder with the adapter.
@@
In Experiment 6, we examined the impact of cross-lingual settings on the performance of ASR representations in NSER. Although ASR systems effectively captured speech and suppressed background noise, their performance degraded when handling languages not included in their training data. Using the CASIA dataset (Mandarin Chinese) as a monolingual scenario, we compared SSL representations under cross-lingual conditions with ASR representations under both monolingual and cross-lingual conditions. For SSL, we evaluated wav2vec 2.0-base11footnotemark: 1, HuBERT-base22footnotemark: 2, and wavLM-base33footnotemark: 3, pre-trained on 960 hours of English Librispeech data, in frozen and fine-tuned configurations. For ASR, we employed two Whisper variants: Whisper-CN444https://huggingface.co/Jingmiao/whisper-small-chinese_base, trained solely on Mandarin data to represent monolingual conditions, and Whisper-EN555https://huggingface.co/openai/whisper-small.en, trained on English data to represent cross-lingual conditions.
@@ IV-C Implementation
Our deep learning models were developed using Python 3.7 and PyTorch 1.11.0. We performed model training and evaluation on a machine equipped with an Intel(R) Xeon(R) Gold 6248 CPU @ 2.50 GHz, 32 GB of RAM, and one NVIDIA Tesla V100 GPU.
In the speech modality, we employed the Whisper-small44footnotemark: 4 as a large-scale ASR model, which is characterized by its architecture comprising 13 layers in both encoder and decoder configurations, each with a hidden layer size of 768. Additionally, we utilized the Conformer-small666https://huggingface.co/speechbrain/asr-conformer-transformerlm-librispeech as the conventional ASR model for this study. To maintain fairness in model sizes, the SSL model utilized in this study is also of the Base size, with a hidden layer size of 768. For the text modality, we employed the BERT-base777https://huggingface.co/google-bert/bert-base-uncased model, featuring a hidden size of 768, 12 attention layers, and 12 attention heads. Consistent with previous studies, the Whisper and Conformer models were frozen during the training process in this study. For the layer adapter module, the embedding size was kept constant at 512 across all layer adapters. For the SER model, each GRU in the architecture consists of two layers with 256 units each and a dropout rate of 0.5. We used the Adam optimizer with a learning rate of 0.0001 and employed the cross-entropy loss function.
| Source | Model | MELD |
| UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | |||||||
| Mel-Spectrogram | SER | 32.73 | 33.41 | 57.12 | 56.10 | 53.71 | 53.34 | 53.22 | 52.61 | |||||
| Conv-TasNet + SER | 33.16 | 33.44 | - | - | 55.03 | 53.84 | 52.21 | 51.68 | ||||||
| DCCRN + SER | 33.39 | 33.24 | - | - | 53.84 | 52.82 | 51.61 | 51.30 | ||||||
| Waveform | Wav2Vec 2.0 + SER | 43.00 | 43.46 | 65.04 | 63.51 | 61.92 | 60.32 | 59.77 | 57.85 | |||||
| Hubert + SER | 42.59 | 42.61 | 66.52 | 65.43 | 61.24 | 60.26 | 60.49 | 60.02 | ||||||
| WavLM + SER | 40.29 | 41.05 | 65.27 | 64.26 | 60.31 | 58.49 | 59.10 | 57.01 | ||||||
| ASR(Conformer) + SER | 39.97 | 40.27 | 72.84 | 73.04 | 68.94 | 69.17 | 66.18 | 65.86 | ||||||
| ASR(Whisper) + SER | 48.95 | 48.87 | 76.50 | 76.53 | 73.00 | 72.42 | 68.79 | 68.39 | ||||||
@@ IV-D Evaluation Metrics
In evaluating our results on IEMOCAP, where a standard train, dev, and test split is not predefined, we employed a five-fold cross-validation approach, consistent with previous studies [49, 48]. For MELD, following established practices in prior research [63, 48], we utilized the Train and Validation sets for training, with the Test set used for evaluation. As for CASIA, we adopted a four-fold cross-validation approach, aligning with previous works [64, 65, 66].
For ASR performance evaluation, we used unweighted average recall (UAR) and F1 scores, which are widely recognized metrics for assessing models on imbalanced datasets [67]. These metrics were applied across four distinct emotional categories. To evaluate the ASR performance, we used word error rate (WER), which is commonly employed to gauge recognition accuracy in speech-related tasks [34].
@@ V-A Effect of ASR Representations on NSER
In this study, we conducted a comprehensive comparative analysis to assess the efficacy of ASR representations in NSER, benchmarking them against conventional noise reduction methods and fine-tuning strategies using SSL models. The synthesized results of this analysis are systematically presented in Table I.
The results highlight the superior performance of ASR representations across various datasets. For the MELD dataset, Whisper representations improve UAR and F1 scores by 16.22 and 15.46 compared with noisy Mel-Spectrograms, and by 15.79 and 15.43 compared with ConvTasNet, the best-performing denoising model in this dataset, respectively. Compared with the best-performing SSL representation (Wav2vec 2.0), Whisper achieves gains of 5.95 in UAR and 5.41 in F1. These results indicate that Whisper effectively captures acoustic and semantic features, outperforming both conventional denoising and SSL approaches. The Conformer model, although effective, scores 8.98 lower in UAR and 8.60 lower in F1 compared with Whisper.
@@
On the IEMOCAP dataset, Whisper outperforms noisy Mel-Spectrograms by 19.38 in UAR and 20.43 in F1. Against the best-performing SSL representation (Hubert), Whisper improves UAR and F1 by 9.98 and 11.10, respectively. Whisper also surpasses the Conformer model by 3.66 in UAR and 3.49 in F1. Similarly, Whisper achieves notable gains on the IEMOCAP Environment subset, improving UAR and F1 by 19.29 and 19.08 compared with noisy Mel-Spectrograms, and by 17.97 and 18.58 over ConvTasNet, respectively. Against the best-performing SSL representation (Wav2vec 2.0), Whisper improves by 11.08 in UAR and 12.10 in F1. Whisper surpasses the Conformer model by 4.06 in UAR and 3.25 in F1.
For the IEMOCAP Human subset, where human noise poses a significant challenge, Whisper shows improvements of 15.57 in UAR and 15.78 in F1 over noisy Mel-Spectrograms. Compared with ConvTasNet, Whisper achieves gains of 16.58 in UAR and 16.71 in F1. Whisper also demonstrates significant advantages over SSL models, with improvements of 8.3 in UAR and 8.37 in F1 compared with the best-performing SSL representation (Hubert). Compared with the Conformer model, Whisper shows additional improvements of 2.61 in UAR and 2.53 in F1.
In summary, ASR representations, particularly Whisper, consistently outperform both conventional denoising and SSL methods across datasets. Whisper demonstrates superior robustness in handling both environmental and speech noise, leveraging rich acoustic and semantic features to achieve significant performance gains. These results underscore its effectiveness for NSER, particularly in challenging noise scenarios.
| Model | MELD |
| UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | |||||||
| Encoder | Last | 44.56 | 45.69 | 71.66 | 70.85 | 66.69 | 66.17 | 61.96 | 61.64 | |||||
| Mean | 46.36 | 46.46 | 72.91 | 72.19 | 68.54 | 67.83 | 65.38 | 65.21 | ||||||
| Adapter | 46.71 | 48.01 | 75.02 | 74.10 | 70.56 | 69.43 | 67.04 | 66.36 | ||||||
| Decoder | Last | 45.55 | 46.55 | 69.83 | 69.41 | 64.90 | 64.06 | 61.80 | 61.13 | |||||
| Mean | 45.00 | 46.58 | 70.70 | 70.29 | 65.92 | 65.62 | 62.07 | 61.42 | ||||||
| Adapter | 46.37 | 47.55 | 72.75 | 72.39 | 69.72 | 68.39 | 66.99 | 66.61 | ||||||
| Encoder + Decoder | Proposed | 48.95 | 48.87 | 76.50 | 76.53 | 73.00 | 72.42 | 68.79 | 68.39 | |||||

### v4

@@ 2.1 IEMOCAP-Human
To replicate human speech noise in realistic acoustic environments, we merged the IEMOCAP data with the Office category from the DEMAND corpus. We used the OMEETING segment of the Office data, recorded in a busy cafeteria, to augment the IEMOCAP dataset.
2.2 IEMOCAP-Environment (Park)
To simulate natural noise in realistic acoustic environments, we combined the IEMOCAP data with the Nature category from the DEMAND corpus. Specifically, we used the NPARK segment of the Nature data, recorded in a bustling park, to augment the IEMOCAP dataset.
2.3 IEMOCAP-Environment (Traffic)
To simulate traffic noise in realistic acoustic environments, we combined the IEMOCAP data with the Transportation category from the DEMAND corpus. We used the Traffic segment of the Transportation data, recorded on a busy roadway, to augment the IEMOCAP dataset.
@@ IV-B Experimental Procedure
In this study, we conducted seven experiments to investigate various aspects of ASR representations in NSER.
In Experiment 1, we evaluated a comprehensive set of baselines for NSER under 0 dB SNR across the MELD, IEMOCAP, IEMOCAP-Human, and IEMOCAP-Environment (Park, Traffic) datasets. As a starting point, we considered the Mel-spectrogram representation as a conventional NSER baseline. We further extracted HSF features using the OpenSMILE toolbox111https://github.com/audeering/opensmile, which provides extensive acoustic feature configurations, including IS09_emotion, IS10_paraling, IS11_speaker_state, IS12_speaker_trait, IS13_ComParE, and eGeMAPS. For noise reduction, we adopted Conv-TasNet for speech separation and DCCRN for speech enhancement. We additionally evaluated mainstream NSER models such as MLDK and MV-BiSSM. In terms of self-supervised learning (SSL) methods, we fine-tuned Wav2Vec 2.0222https://huggingface.co/facebook/wav2vec2-base, HuBERT333https://huggingface.co/facebook/hubert-base-ls960, and WavLM444https://huggingface.co/microsoft/wavlm-base on noisy speech. Finally, we compared ASR-based representations, including a Conformer-based ASR model as the conventional baseline [22, 66, 60], as well as large-scale ASR models such as SpeechLM [80] and Whisper, all integrated with the same adapter mechanism on both encoder and decoder layers as in our proposed method. A summary of all baselines is provided in Table I.
| Method | Model | Description |
| – | BiGRU | Traditional NSER baseline using Mel-Spectrogram. |
| HSF | IS09_emotion [57] | 384 features; statistical functionals of basic Low-Level Descriptors (LLDs, e.g., energy, pitch). |
| IS10_paraling [54] | 1582 features; 34 LLDs + deltas with functionals, plus pitch features. | |
| IS11_speaker_state [53] | 4368 features; multiple LLD groups with utterance-level functionals. | |
| IS12_speaker_trait [55] | 6125 features; extends IS11 with local-extrema functionals. | |
| IS13_ComParE [56] | 6373 features; comprehensive set of energy, spectral, cepstral, and voicing descriptors. | |
| eGeMAPS [15] | 88 features; compact, theory-driven set with MFCCs, spectral flux, and voiced/unvoiced statistics. | |
| Denoising | Conv-TasNet [41] | Convolutional time-domain network for speech separation. |
| DCCRN [28] | Deep complex convolutional recurrent network for speech enhancement. | |
| NSER | MLDK [39] | Multi-level deep kernel learning model for NSER. |
| MV-BiSSM [38] | Multi-view bidirectional sequential self-attention model for NSER. | |
| SSL | Wav2Vec 2.0 [3] | Self-supervised model for learning contextualized speech representations. |
| HuBERT [27] | Self-supervised model for learning robust speech embeddings through hidden-unit clustering. | |
| WavLM [9] | Self-supervised model for learning speaker- and environment-aware speech representations. | |
| ASR | Conformer [22] | End-to-end ASR backbone combining convolution and transformer layers to effectively capture both local dependencies (e.g., phonetic details) and global context for speech recognition. |
| SpeechLM [80] | Multilingual speech–text pre-trained ASR model that unifies acoustic and linguistic information through joint objectives, enabling stronger transfer across languages and tasks. | |
| Whisper [49] | Large-scale encoder–decoder ASR model designed for robust multilingual and multitask speech recognition, covering diverse languages, accents, and noisy acoustic conditions. |
In Experiment 2, we examined the effectiveness of different ASR representations for NSER using the MELD, IEMOCAP, IEMOCAP-Human, and IEMOCAP-Environment (Park, Traffic) datasets under 0 dB SNR conditions. Specifically, we compared: (i) single-layer representations (either the last layer or the best-performing layer among all layers), and (ii) fusion-based multi-layer representations, which were implemented through three strategies: simple fusion (mean pooling across layers), parameterized fusion (scalar mixing and weighted sum), and advanced fusion (attention-based pooling). In addition, we proposed the Layer Adapter approach, which extends weighted-sum fusion of ASR encoder and decoder representations by enabling more effective cross-layer integration, thereby producing richer and more robust features for NSER.
In Experiment 3, we examined the layer-wise effectiveness of ASR representations for NSER. Specifically, we compared representations extracted from the encoder and decoder layers of ASR models, whereas for SSL models only encoder-layer representations were considered. All experiments were conducted on the IEMOCAP, IEMOCAP-Human, and IEMOCAP-Environment (Park, Traffic) datasets at an SNR of 0 dB.
@@
In Experiment 4, we examined the robustness of ASR representations for NSER under varying noise intensities. We conducted a comparative analysis using the IEMOCAP-Human and IEMOCAP-Environment (Park, Traffic) datasets, systematically controlling noise at SNR levels of -10, -5, 0, 5, and 10 dB. We evaluated different ASR representations, including: Whisper encoder with the Layer Adapter; Whisper decoder with the Layer Adapter; Combined Whisper encoder and decoder with the Layer Adapter, alongside representations from SSL models.
@@
In Experiment 5, we investigated the correlation between ASR performance and NSER outcomes across different noise intensities. The comparative analysis was conducted using the IEMOCAP-Human dataset and the IEMOCAP-Environment dataset (Park, Traffic). For evaluation, we considered ASR representations from two models: a Conformer-based ASR model as a conventional baseline, and Whisper, a large-scale ASR model.
@@
In Experiment 6, we evaluated the effectiveness of ASR representations across different modalities for NSER, using the MELD, IEMOCAP, IEMOCAP-Human, and IEMOCAP-Environment (Park, Traffic) datasets under 0 dB SNR. We conducted a comparative analysis between speech- and text-based modalities, leveraging SSL methods on both ASR-generated and ground-truth transcriptions. For the text modality, we employed a BERT model, following prior studies on emotion recognition [1, 32]. For the speech modality, we compared ASR representations from Whisper, including the encoder with an adapter, the decoder with an adapter, and a combined encoder–decoder configuration with an adapter.
@@
In Experiment 7, we examined the impact of cross-lingual settings on the performance of ASR representations in NSER. Although ASR systems effectively captured speech and suppressed background noise, their performance degraded when handling languages not included in their training data. Using the CASIA dataset (Mandarin Chinese) as a monolingual scenario, we compared SSL representations under cross-lingual conditions with ASR representations under both monolingual and cross-lingual conditions. For SSL, we evaluated wav2vec 2.0-base11footnotemark: 1, HuBERT-base22footnotemark: 2, and WavLM-base33footnotemark: 3, pre-trained on 960 hours of English Librispeech data, in frozen and fine-tuned configurations. For ASR, we employed two Whisper variants: Whisper-CN555https://huggingface.co/Jingmiao/whisper-small-chinese_base, trained solely on Mandarin data to represent monolingual conditions, and Whisper-EN666https://huggingface.co/openai/whisper-small.en, trained on English data to represent cross-lingual conditions. Moreover, since CASIA provides parallel utterances with consistent textual content across different emotions, this experiment mitigates the confounding influence of linguistic information. Consequently, the results more faithfully reflect the contributions of paralinguistic cues captured by ASR representations in NSER.
@@ IV-C Implementation
All models were implemented in Python 3.7 with PyTorch 1.11.0. Training and evaluation were conducted on a server equipped with an Intel(R) Xeon(R) Gold 6248 CPU @ 2.50 GHz, 32 GB RAM, and a single NVIDIA Tesla V100 GPU.
In the speech modality, we employed Conformer-small777https://huggingface.co/speechbrain/asr-conformer-transformerlm-librispeech as a representative conventional ASR model, and additionally considered two large-scale approaches: (i) Whisper-small44footnotemark: 4 and (ii) SpeechLM-Base888https://github.com/microsoft/SpeechT5/tree/main/SpeechLM , both adopting a transformer encoder–decoder architecture with 12 encoder layers, 12 decoder layers, and a hidden size of 768. For the text modality, we used BERT-base999https://huggingface.co/google-bert/bert-base-uncased , featuring a hidden size of 768, 12 attention layers, and 12 heads.
For baseline experiments, we implemented several comparison models. For Mel-spectrogram and denoised Mel-spectrogram features, a BiGRU with 256 units and dropout 0.5 was employed. For HSF features, we adopted a fully connected network with 512 units and dropout 0.5. For SSL-based methods, mean pooling was applied along the temporal dimension, followed by a fully connected classifier. To ensure fairness in model capacity, all SSL models were of Base size (hidden size 768). Detailed settings are shown in Table II.
| Baseline | Mel / Denoising | HSF | SSL |
| Hyperparameters | |||
| Batch size | 32 | 32 | 32 |
| Epochs | 50 | 50 | 50 |
| Learning rate | |||
| Dropout | 0.5 | 0.5 | 0.5 |
| Optimizer | Adam | Adam | Adam |
| Loss | Cross-Entropy | Cross-Entropy | Cross-Entropy |
| Model configurations | |||
| Backbone frozen | No | No | No |
| Hidden size | 256 | 512 | 768 |
| Downstream head | BiGRU | FC | MeanPool + FC |
For ASR-based models, consistent with prior work, the Whisper, Conformer, and SpeechLM backbones were frozen during training. The proposed Layer Adapter employed a BiGRU with 256 units per direction and a dropout rate of 0.5, with the embedding size fixed at 512 across adapters. The emotion classifier was a fully connected layer with a dropout of 0.5. Training used a batch size of 32, the Adam optimizer with a learning rate of , and cross-entropy loss.
@@ IV-D Evaluation Metrics
In evaluating our results on IEMOCAP, where a standard train, dev, and test split is not predefined, we employed a five-fold cross-validation approach, consistent with previous studies [64, 61]. For MELD, following established practices in prior research [20, 61], we utilized the Train and Validation sets for training, with the Test set used for evaluation. As for CASIA, we adopted a four-fold cross-validation approach, aligning with previous work [47].
For NSER performance evaluation, we used unweighted average recall (UAR) and F1 scores, which are widely recognized metrics for assessing models on imbalanced datasets [8]. These metrics were applied across four distinct emotional categories. To evaluate the ASR performance, we used word error rate (WER), which is commonly employed to gauge recognition accuracy in speech-related tasks [67].
@@ V-A Effect of ASR Representations on NSER
In this study, we conducted a comprehensive comparative analysis to assess the efficacy of ASR representations in NSER, benchmarking them against traditional HSF features, noise reduction approaches, mainstream NSER models, and SSL methods. The results are presented in Table III.
| Method | Feature | MELD | IEMOCAP | IEMOCAP (Human) | IEMOCAP (Environment) | ||||||
| Park | Traffic | ||||||||||
| UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | ||
| – | Mel-Spectrogram | 32.73 | 33.41 | 57.12 | 56.10 | 53.22 | 52.61 | 53.71 | 53.34 | 54.00 | 52.83 |
| HSF | IS09_emotion | 31.92 | 31.28 | 58.65 | 57.84 | 56.31 | 55.43 | 54.73 | 54.33 | 54.81 | 53.87 |
| IS10_paraling | 34.90 | 35.34 | 61.22 | 59.49 | 56.34 | 55.96 | 56.15 | 55.01 | 56.31 | 55.43 | |
| IS11_speaker_state | 37.06 | 35.88 | 60.17 | 59.83 | 55.12 | 54.06 | 55.57 | 53.96 | 55.12 | 54.06 | |
| IS12_speaker_trait | 35.22 | 35.10 | 60.46 | 59.11 | 55.31 | 53.56 | 54.38 | 52.80 | 55.31 | 53.56 | |
| IS13_ComParE | 37.33 | 36.04 | 60.11 | 59.30 | 54.84 | 53.31 | 55.85 | 54.42 | 54.84 | 53.31 | |
| eGeMAPS | 34.04 | 33.36 | 58.79 | 58.02 | 53.47 | 52.46 | 54.75 | 54.02 | 53.95 | 53.41 | |
| Denoising | Conv-TasNet | 33.16 | 33.44 | - | - | 52.21 | 51.68 | 55.03 | 53.84 | 55.62 | 54.17 |
| DCCRN | 33.39 | 33.24 | - | - | 51.61 | 51.30 | 53.84 | 52.82 | 55.12 | 53.48 | |
| NSER | MLDK | 45.51 | 45.25 | 70.65 | 69.11 | 65.37 | 63.58 | 68.82 | 67.29 | 70.21 | 68.88 |
| MV-BiSSM | 47.71 | 47.45 | 75.89 | 74.30 | 68.51 | 66.74 | 72.66 | 72.01 | 73.97 | 73.52 | |
| SSL | Wav2Vec 2.0 (FT) | 43.00 | 43.46 | 68.08 | 67.19 | 59.24 | 57.49 | 60.68 | 59.52 | 64.12 | 62.63 |
| HuBERT (FT) | 42.59 | 42.61 | 66.91 | 66.32 | 61.08 | 60.08 | 61.41 | 60.72 | 62.66 | 62.04 | |
| WavLM (FT) | 40.29 | 41.05 | 65.29 | 64.22 | 59.28 | 57.36 | 60.97 | 58.74 | 62.28 | 60.95 | |
| ASR | Conformer (FZ) | 39.97 | 40.27 | 72.84 | 73.04 | 66.18 | 65.86 | 68.94 | 69.17 | 70.54 | 70.96 |
| SpeechLM (FZ) | 44.27 | 45.68 | 73.18 | 74.79 | 66.59 | 66.74 | 69.06 | 70.49 | 71.71 | 71.14 | |
| Proposed | 48.95 | 48.87 | 76.50 | 76.53 | 68.79 | 68.39 | 73.00 | 72.42 | 74.39 | 74.04 | |
The results highlight the superior performance of ASR representations across various datasets. On the MELD dataset, the proposed method achieves the largest gains. Compared with noisy Mel-Spectrograms, it improves by 16.22% in UAR and 15.46% in F1. Against the best HSF features (IS13), the gains are 11.62% and 12.83%. Over denoising baselines, the proposed method achieves 15.56% improvement in UAR and 15.63% in F1 compared with DCCRN. Relative to the strongest NSER baseline (MV-BiSSM), it adds 1.24% in UAR and 1.42% in F1. Compared with Wav2Vec 2.0, the best SSL representation, the proposed approach achieves 5.95% and 5.41% improvements. Finally, even against SpeechLM, a large-scale ASR model, the proposed model still surpasses by 4.68% in UAR and 3.19% in F1.
On the IEMOCAP dataset, the proposed approach improves by 19.38% in UAR and 20.43% in F1 over noisy Mel-Spectrograms. Relative to the HSF feature (IS10), it gains 15.28% in UAR and 17.04% in F1. Against MV-BiSSM, the strongest NSER baseline, the proposed method provides 0.61% and 2.23% improvements. Compared with Wav2Vec 2.0, the best SSL, the gains are 8.42% and 9.34%. Over SpeechLM, it improves by 3.32% in UAR and 1.74% in F1.
For the IEMOCAP-Human subset, the proposed method shows 15.57% and 15.78% improvements over noisy Mel-Spectrograms. Compared with the best HSF features (IS10), the gains reach 12.45% in UAR and 12.43% in F1. Over Conv-TasNet, it improves by 16.58% and 16.71%. Compared with MV-BiSSM, the proposed method is slightly higher, achieving 0.28% and 1.65% gains. Against HuBERT, the best SSL, it achieves 7.71% and 8.31% gains, and over SpeechLM it further improves by 2.20% and 1.65%.
@@
For the IEMOCAP-Environment subset (Park), the proposed method improves by 19.29% and 19.08% compared with noisy Mel-Spectrograms. Over the best HSF features (IS10), it achieves 16.85% and 17.41% improvements. Against ConvTasNet, the gains are 17.97% and 18.58%. Compared with MV-BiSSM, the proposed method is slightly higher, with improvements of 0.34% in UAR and 0.41% in F1. Compared with HuBERT, the best SSL, it improves by 11.59% and 11.70%. Over SpeechLM, the gains are 3.94% in UAR and 1.93% in F1.
For the IEMOCAP-Environment subset (Traffic), the proposed method shows improvements of 20.39% and 21.21% over noisy Mel-Spectrograms. Compared with the best HSF features (IS10), the gains are 18.08% in UAR and 18.61% in F1. Over Conv-TasNet, it improves by 18.77% and 19.87%. Relative to MV-BiSSM, the proposed method improvements of 0.42% in UAR and 0.52% in F1. Compared with Wav2Vec 2.0, the best SSL model, it achieves 10.27% and 11.41% improvements. Over SpeechLM, the gains are 2.68% in UAR and 2.90% in F1.
In summary, the proposed ASR-based approach, exemplified by the Whisper encoder–decoder framework, achieves consistently strong performance over HSF features, conventional denoising, prior NSER baselines, and SSL models, while demonstrating notable robustness to environmental and human noise through joint use of encoder and decoder representations.